OpenStudy (anonymous):

A water drum in the shape of a rectangular prism measures 5m by 4m by 2m. It is completely full of water. The water is siphoned into a cylinder tank of diameter 3m. How far up the side of the tank will the water reach?

OpenStudy (jhannybean):



OpenStudy (jhannybean):

Hmm...volume of cylindrical tub would be... \[\large V= \pi r^2 h\] and i believe you're solving for h. You're going to create a ratio for the Volume of the rectangular prism to the volumeof the cylindrical tub to find h.

OpenStudy (anonymous):

no, \[\Pi \times r ^{2} \times h\] here r= 3/2=1.5 everythin else is correct